OpenStudy (anonymous):

Use L'Hospital Rule limx->infinity x^3*e^-x^2

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} x^{3}*e^{-x^2}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} x^3*e^{-x^2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just do not know how to type it in that way when writing the original question

OpenStudy (agent0smith):

You could differentiate a couple times, since it's indeterminate, and eventually x^3 becomes 0.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{ x^3 }{ e^{x^2} }\]

OpenStudy (agent0smith):

^ like that. Just from glancing at it, it'll approach zero (e^x approaches 0 faster than x^3 approaches infinity) - this is a good way to gauge your answer.

OpenStudy (anonymous):

how do you know it becomes 0. I thought it would go to infinity becasue I can not get rid of the e^x

OpenStudy (agent0smith):

exponential functions grow at a faster rate than polynomials. Or in this case, shrink at a faster rate.

OpenStudy (anonymous):

yes, i got that first step asaad123

OpenStudy (agent0smith):

Now use l hopital and differentiate the numerator and denominator.

OpenStudy (anonymous):

ooooo, soo smart. geeze, ty

OpenStudy (agent0smith):

So you can often know the answer before proving it, which helps to check your answer. affter differenitating once....\[\Large \lim_{x \rightarrow \infty}\frac{ 3x^2 }{ 2x e^{x^2} }\]

OpenStudy (anonymous):

so basiaclly because i would set it up for the exp to go in the bottom itll always go to zero regardless of what the num is right?

OpenStudy (anonymous):

because nothing grows faster than an exp

OpenStudy (agent0smith):

Provided your a in a^x is bigger than 1 If it's something like a polynomial over an exponential, yes \[\Large \frac{ x^9 }{ 1.1^x}\] eventally this should approach zero.

OpenStudy (agent0smith):

And if you graph it, it does... eventually. https://www.google.com/search?q=x%5E9%2F1.1%5Ex&oq=x%5E9%2F1.1%5Ex&aqs=chrome..69i57.5734j0&sourceid=chrome&espvd=210&es_sm=93&ie=UTF-8 zoom out a lot and you'll see it.

OpenStudy (anonymous):

so if i was to keep d/dx it ill eventually get 3/infin right?

OpenStudy (anonymous):

which =0

OpenStudy (anonymous):

right

OpenStudy (anonymous):

TYsoVM :)

OpenStudy (agent0smith):

\[\Large \lim_{x \rightarrow \infty}\frac{ 3x^2 }{ 2x e^{x^2} }\] simplify it first, \[\Large \lim_{x \rightarrow \infty}\frac{ 3x }{ 2 e^{x^2} }\] now differntiate again.

OpenStudy (anonymous):

gotcha